# Circumbinary disks in post common envelope binary systems with compact objects

### v1

@@ Abstract
We conduct a population synthesis study using the binary population synthesis code compas to explore the formation of circumbinary disks (CBDs) following the common envelope evolution (CEE) phase of a giant star and a neutron star (NS) or black hole (BH). We focus on massive binary systems that evolve into double compact object (DCO) binaries after the exposed core of the giant collapses to form a second NS or BH. A CBD around the binary system of the giant’s core and the compact object lives for a short time at the termination of the CEE phase and alters the orbital evolution of the binary. We parameterize the conditions for CBD formation in post-CEE binaries and present characteristics of DCO progenitors that are likely or unlikely to form CBDs. We find that CBD formation is most common in BH-BH binaries and NS-NS binaries that are expected to merge within Hubble time. Furthermore, we find that the interaction of the CBD with the core - NS/BH system at the termination of the CEE reduces the expected rate of DCO mergers, regardless of whether these binaries tighten or expand due to this interaction. If the binary system loses angular momentum to the CBD, it may produce a luminous transient due to a merger between the NS/BH and the core of the giant rather than gravitational wave sources. Thus, accounting for post-CEE CBD formation and its interaction with the binary system in population synthesis studies is significant for obtaining reliable predictions of the gravitational wave event rates expected by current detectors.
@@ 1 INTRODUCTION
The common envelope evolution (CEE) phase involves a wide range of processes on various timescales and length-scales, many of which are very difficult to simulate even with the most sophisticated existing hydrodynamical numerical codes (for a review see, e.g., Röpke & De Marco 2023). Three-dimensional hydrodynamical simulations of CEE over four decades (e.g., de Kool 1987; Livio & Soker 1988; Terman et al. 1994; Rasio & Livio 1996; Sandquist et al. 1998; Ricker & Taam 2008; Passy et al. 2012; Nandez et al. 2014; Ivanova & Nandez 2016; Ohlmann et al. 2016; Staff et al. 2016; Iaconi et al. 2017; Prust & Chang 2019; Sand et al. 2020; Glanz & Perets 2021a, b; Lau et al. 2022; Hillel et al. 2023; Schreier et al. 2023; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier & Pejcha 2024a; González-Bolívar et al. 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024) and of the grazing envelope evolution (e.g., Shiber et al. 2017, 2019; Shiber & Iaconi 2024), have had difficulties including all the CEE sub-phases together and all physical processes required to fully simulate the CEE phase.
@@
One process the simulations can not fully follow is the possible formation of a circumbinary disk (CBD) around the surviving binary system. If the dynamical plunge-in of the accretor is stalled, the remnant binary could interact with the flattened envelope outside the binary orbit over a much longer timescale, resulting in the formation of a CBD (e.g., Kashi & Soker 2011). The re-distribution of angular momentum and energy in the envelope gas leads to a torus-like structure around the core of the primary and the secondary stars. Current three-dimensional hydrodynamical simulations are too resource-demanding to follow the interaction of the remnant binary system with the envelope over multiple dynamical timescales as needed for CBD formation. Hydrodynamical simulations of the CEE, however, form a flattened envelope (e.g., Sandquist et al. 1998) with a funnel along the symmetry axis (e.g., Chamandy et al. 2018, 2020; Zou et al. 2020, 2022; Ondratschek et al. 2022; Vetter et al. 2024) as predicted by an analytical calculation many years earlier (Soker, 1992), showing the potential of CBD formation.
@@
The interaction of the post-CEE CBD with the binary system was suggested to influence the subsequent binary evolution (e.g., Kashi & Soker 2011; Chen & Podsiadlowski 2017; Muñoz et al. 2019; D’Orazio & Duffell 2021; Tuna & Metzger 2023; Gagnier & Pejcha 2023; Wei et al. 2024; Siwek et al. 2023; Valli et al. 2024; Vetter et al. 2024). Hydrodynamical simulations that aim to follow the evolution of CBDs and their long-term interaction with the binary system they encompass encounter multiple numerical challenges. For instance, properly resolving the cavity that contains the binary is crucial for estimating the torques that dictate the orbital evolution (e.g., Gagnier & Pejcha 2023). Since the core of the giant star and its companion are usually represented as point masses with artificial gravitational potentials, the gravitational torques are inherently inaccurate, leading to imprecise orbital separations. Another layer of complexity in simulating the CBD-binary interaction is posed by correctly modeling magnetic fields in the post-dynamical in-spiral phase of CEE (e.g., Gagnier & Pejcha 2024b). Although magnetic fields are thought to have negligible impact on binary orbital separations, their amplification could lead to outflows that alter the morphology and densities around the disk, influencing the orbital evolution.
@@
In post-common envelope systems of massive stars, the interaction of the CBD with the binary could lead to detectable transients. If the binary system contains a neutron star (NS) or a black hole (BH) and the stripped core of a giant, then mass accretion by the compact object from the CBD could result in the launching of jets that power a luminous event. Tuna & Metzger (2023) suggest that jet-powered radio emission from such a system might be compatible with expectations from fast blue optical transients. In particular, long-lived disks could result in delayed explosions with less hydrogen as in the case of AT2018cow-like transients or type Ibn/Icn supernovae (e.g., Soker 2022; Metzger 2022; Tuna & Metzger 2023; Wei et al. 2024). Furthermore, dust formation by the jets launched from the CBD could result in detectable IR radiation (see Grichener 2024 for a discussion on potential observational signatures of such systems). In the case of much wider post-asymptotic giant branch binaries, having orbital periods of $\simeq 100-1000$ days and masses of $M_{1}\simeq 1M_{\odot}$, mass accretion and jets launching from a CBD of several AU could shape a bipolar nebula (e.g., Bollen et al. 2022; Verhamme et al. 2024). These CBDs, however, are expected to be post-grazing envelope evolution disks rather than CEE (e.g., Soker 2020).
@@
One approach to allow a wide exploration of CEE and post-CEE outcomes is parameterizing processes. This is particularly critical in performing population synthesis studies. We aim to incorporate the possibility of forming a post-CEE CBD into population synthesis; here, we re-analyze the data from Grichener (2023) (publicly available at doi:10.5281/zenodo.11237179) generated by the rapid population synthesis code compas (Riley et al., 2022), as described in Section 2. Towards our goal, in Section 3, we build a toy model to parameterize the conditions to form post-CEE CBDs. In Section 4 we employ our toy model to the population synthesis data of massive binary systems that end with either NS - NS, BH - BH, or NS - BH binary systems. In Section 5, we discuss the potential implications of our results for the formation and merger rates of double compact object (DCO) binaries. We summarize and discuss the effect of extending the parameter space on our results in Section 6.
2 Methods
@@
We use rapid population synthesis to follow the evolution of massive binaries and find the properties of systems where a circumbinary disk would form in the post-CEE phase of a compact object (NS or BH) and a giant star. We analyze datasets from Grichener (2023) [^1] Publicly available at doi:10.5281/zenodo.11237179. generated by version 02.31.06 of the population synthesis code compas (Compact Object Mergers: Population Astrophysics and Statistics; Stevenson et al. 2017; Vigna-Gómez et al. 2018; Riley et al. 2022).
@@
Given initial distributions and parametric assumptions, compas evolves isolated binary systems using the analytical fits from Hurley et al. (2000) and Hurley et al. (2002). The mass distribution of the primary (initially more massive) star is determined by the Kroupa initial mass function in the form $dN/dM_{\rm 1,ZAMS}\propto M_{\rm 1,ZAMS}^{-2.3}$ (Kroupa, 2001). The range of the masses $5\leq M_{\rm 1,ZAMS}/M_{\odot}\leq 100$ was chosen to account for primaries that could result in a supernova explosion and the formation of a compact object. [^2] The stars with the lowest masses in our distribution can gain mass through mass transfer and become massive enough to explode as supernovae. The mass of the secondary (initially less massive) star was determined by a flat distribution for a mass ratio between $0.1\leq q_{\text{ZAMS}}\equiv M_{2,\text{ZAMS}}/M_{1,\text{ZAMS}}\leq 1$ and the semi-major axis was determined by a flat-in-log distribution in the range $0.1\leq a_{\rm ZAMS}\rm/AU\leq 1000$ (Sana et al., 2012). For the NS natal-kicks induced by supernova explosions we take a bimodal velocity distribution with a higher mode of $\sigma_{\rm high}=265~{}\rm{km}~{}\rm{s}^{-1}$ for regular core collapse supernovae and a lower mode of $\sigma_{\rm high}=30~{}\rm{km}~{}\rm{s}^{-1}$ corresponding to electron capture supernovae and ultra stripped supernovae, while for the BHs in our sample we reduce this natal-kick distribution by a fallback factor. We focus on sub-populations of $10^{7}$ binaries with solar metallicity ($Z_{\rm\odot}=0.0142$; Asplund et al. 2009) and common envelope efficiency parameter $\alpha_{\rm CE}=1$ (equation 1) as in the ’fiducial’ model of Vigna-Gómez et al. (2018).
@@
In some of the systems sampled in Grichener (2023) there were episodes of dynamically unstable mass transfer that led to CEE. compas determines the stability of mass transfer by comparing the donor’s response to mass loss with that of the Roche-lobe radius: if the donor’s radius response, $d\ln R_{\ast,i}/d\ln M_{\ast,i}$, where $R_{\ast,i}$ and $M_{\ast,i}$ are the radius and mass of the donor at the onset of mass transfer, is greater than or equal to that of the Roche lobe, $d\ln R_{\mathrm{RL}}/d\ln M_{\ast,i}$, where $R_{\mathrm{RL}}$ is the Roche radius, the mass transfer is stable. Otherwise, it results in a CEE event (e.g., Paczyński & Sienkiewicz 1972). To find the value of the semi-major axis immediately after the CEE phase, compas uses the energy formalism  [^3] Other methods use angular momentum parametrization to estimate the final orbital separation (e.g., Nelemans & Tout 2005; Toonen et al. 2012; Di Stefano et al. 2023). However, the angular momentum prescriptions could be problematic (e.g., Webbink 2008; Cohen & Soker 2023).. In this formalism, we compare the binding energy of the envelope $E_{\rm bind}$ to the orbital energy before and after the CEE phase (e.g., van den Heuvel 1976; Iben & Tutukov 1984; Livio & Soker 1988; Ivanova et al. 2013; Iaconi & De Marco 2019)
@@
where $\alpha_{\rm CE}$ is the common envelope efficiency parameter that determines which fraction of the orbital energy will go to unbinding the envelope, $M_{\rm 1,i}$ and $M_{\rm 2,i}$ are the masses of the primary and secondary stars at the onset of CEE, respectively, $M_{\rm 1,f}$ and $M_{\rm 2,f}$ are the masses of the primary and secondary stars after CEE, respectively, $a_{\rm i}$ is the semi-major axis of the binary systems before the CEE phase and $a_{\rm f}$ is the semi-major axis of the binary systems at the end of CEE. For more information about the population synthesis model assumptions and initial distributions see Section 2 in Grichener (2023).
@@
We focus on sub-populations in which a compact object enters a CEE phase with a giant star, where the formation of a CBD after CEE could potentially lead to shrinkage in the orbital separation and a post-CEE merger (e.g., Tuna & Metzger 2023; Wei et al. 2024), accounting for various luminous transients (e.g., Grichener 2024). Without taking the formation and interaction of the binary with the CBD into consideration, most of these systems end as DCO binaries, implying a potential reduction in their formation and merger rates (see Section 5). compas allows us to examine which parameters and stellar quantities can affect the process of creating a CBD. To find the properties of system that will form a post-common envelope CBD, we compare the specific angular momentum of the envelope at the end of the CEE phase to the Keplerian specific angular momentum around the binary system, as we further elaborate in Section 3.
3 Parametrization of post-common envelope circumbinary disk formation
3.1 The toy model
@@
We build a toy model to parameterize the conditions for forming a post-CEE CBD and incorporating it in population synthesis studies. We estimate the fraction of the envelope angular momentum at the beginning of the CEE that the leftover envelope maintains and equate it to the minimum angular momentum that the envelope requires to form a CBD. We then implement this parametrization in the data of Grichener (2023) and present our results in Sections 4 and 5.
@@
We assume that at the entrance to the CEE phase, the secondary star, an NS or a BH in our study, manages to spin up the envelope to synchronization. We take the envelope to have a solid-body rotation with the break-up angular velocity $\omega_{\rm bu}=\sqrt{GM_{\ast,i}/R^{3}_{\ast,i}}$, where $M_{\ast,i}$ and $R_{\ast,i}$ are the mass and radius of the red supergiant star at the onset of CEE. For the moment of inertia of the envelope at this stage, we take $I_{\rm env,i}=k_{\rm e}M_{\rm env,i}R^{2}_{\rm\ast,i}$, where $M_{\rm env,i}$ is the envelope mass at the onset of CEE, and $k_{\rm e}\simeq 0.2$ for extended envelopes of giants (e.g., Soker 2004).
We take the post-CEE specific angular momentum of the envelope to be a fraction $\beta$ of the specific envelope angular momentum at the onset of the CEE phase
$j_{\rm env,f}=\beta j_{\rm env,i}=\beta\frac{I_{\rm env,i}\omega_{\rm bu}}{M_{\rm env,i}}=\beta k_{\rm e}R^{2}_{\rm\ast,i}\omega_{\rm bu}.$ (2)
The condition to form a centrifugally-supported CBD is that the specific angular momentum of the material in the CBD is larger than the Keplerian specific angular momentum around the binary system
$j_{\rm env,f}\gtrsim\sqrt{G(M_{\rm core}+M_{\rm CO})a_{\rm f}},$ (3)
where $M_{\rm core}$ is the mass of the giant’s core that becomes stripped after CEE, and $M_{\rm CO}$ is the mass of the compact object companion. From equations (2) and (3) combined with the expression for $\omega_{\rm bu}$, we find the condition to form a CBD from the leftover envelope to be
$\eta_{\rm SIF}\equiv\frac{a_{\rm f}}{R_{\ast,i}}\lesssim(\beta k_{\rm e})^{2}\frac{M_{\ast,i}}{M_{\rm core}+M_{\rm CO}},$ (4)
where $\eta_{\rm SIF}$ is the spiral-in factor of the compact object. The parameter $\beta$ includes also the uncertainties in equation (3), like the exact inner radius of the CBD and the deviation from synchronization at the onset of CEE.
3.2 The role of the parameter $\beta$

### v2

@@ Abstract
We conduct a population synthesis study using the binary population synthesis code compas to explore the formation of circumbinary disks (CBDs) following the common envelope evolution (CEE) phase of a giant star and a neutron star (NS) or black hole (BH). We focus on massive binary systems that evolve into double compact object (DCO) binaries after the exposed core of the giant collapses to form a second NS or BH. A CBD around the binary system of the giant’s core and the compact object alters the orbital evolution of the binary. We parameterize the conditions for CBD formation in post-CEE binaries and present characteristics of DCO progenitors that are likely or unlikely to form CBDs. We find that CBD formation is most common in BH-BH binaries and NS-NS binaries that are expected to merge within Hubble time. Furthermore, we find that the interaction of the CBD with the core - NS/BH system at the termination of the CEE reduces the expected rate of DCO mergers, regardless of whether these binaries tighten or expand due to this interaction. If the binary system loses angular momentum to the CBD, it may produce a luminous transient due to a merger between the NS/BH and the core of the giant rather than gravitational wave sources. Thus, accounting for post-CEE CBD formation and its interaction with the binary system in population synthesis studies is significant for obtaining reliable predictions of the gravitational wave event rates expected by current detectors.
@@ I INTRODUCTION
The common envelope evolution (CEE) phase involves a wide range of processes on various timescales and length-scales, many of which are very difficult to simulate even with the most sophisticated existing hydrodynamical numerical codes (for a review see, e.g., Röpke and De Marco 2023). Three-dimensional hydrodynamical simulations of CEE over four decades (e.g., de Kool 1987; Livio and Soker 1988; Terman et al. 1994; Rasio and Livio 1996; Sandquist et al. 1998; Ricker and Taam 2008; Passy et al. 2012; Nandez et al. 2014; Ivanova and Nandez 2016; Ohlmann et al. 2016; Staff et al. 2016; Iaconi et al. 2017; Prust and Chang 2019; Sand et al. 2020; Glanz and Perets 2021b, a; Lau et al. 2022; Hillel et al. 2023; Schreier et al. 2023; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier and Pejcha 2024a; González-Bolívar et al. 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024) and of the grazing envelope evolution (e.g., Shiber et al. 2017, 2019; Shiber and Iaconi 2024), have had difficulties including all the CEE sub-phases together and all physical processes required to fully simulate the CEE phase.
@@
One process the simulations have difficulties following is the possible formation of a circumbinary disk (CBD) around the surviving binary system. If the dynamical plunge-in of the accretor is stalled, the remnant binary could interact with the flattened envelope outside the binary orbit over a much longer timescale, resulting in the formation of a CBD (e.g., Kashi and Soker 2011). The re-distribution of angular momentum and energy in the envelope gas leads to a torus-like structure around the core of the primary and the secondary stars. Current three-dimensional hydrodynamical simulations of the CEE face challenges in capturing this process in detail. While many of them find that after the ejection of the envelope its gas remains nearby forming a flattened envelope (e.g., Sandquist et al. 1998) or a collimated bipolar structure along the symmetry axis (e.g., Chamandy et al. 2018, 2020; Zou et al. 2020, 2022; Ondratschek et al. 2022) as previously predicted by an analytical calculation (Soker, 1992), showing the potential of CBD formation, simulations resolving the torus-like structure (e.g., Vetter et al. 2024) are scarce.
@@
The interaction of the post-CEE CBD with the binary system was suggested to influence the subsequent binary evolution (e.g., Kashi and Soker 2011; Chen and Podsiadlowski 2017; Muñoz et al. 2019; D’Orazio and Duffell 2021; Tuna and Metzger 2023; Gagnier and Pejcha 2023; Wei et al. 2024; Siwek et al. 2023; Valli et al. 2024; Vetter et al. 2024). Hydrodynamical simulations that aim to follow the evolution of CBDs and their long-term interaction with the binary system they encompass encounter multiple numerical challenges. For instance, properly resolving the cavity that contains the binary is crucial for estimating the torques that dictate the orbital evolution (e.g., Gagnier and Pejcha 2023). Since the core of the giant star and its companion are usually represented as point masses with artificial gravitational potentials, the gravitational torques are inherently inaccurate, leading to imprecise orbital separations. Another layer of complexity in simulating the CBD-binary interaction is posed by correctly modeling magnetic fields in the post-dynamical in-spiral phase of CEE (e.g., Gagnier and Pejcha 2024b). Although magnetic fields are thought to have negligible impact on binary orbital separations, their amplification could lead to outflows that alter the morphology and densities around the disk, influencing the orbital evolution.
@@
In post-common envelope systems of massive stars, the interaction of the CBD with the binary could lead to detectable transients. If the binary system contains a neutron star (NS) or a black hole (BH) and the stripped core of a giant, then mass accretion by the compact object from the CBD could result in the launching of jets that power a luminous event. Tuna and Metzger (2023) suggest that jet-powered radio emission from such a system might be compatible with expectations from fast blue optical transients. In particular, long-lived disks could result in delayed explosions with less hydrogen as in the case of AT2018cow-like transients or type Ibn/Icn supernovae (e.g., Soker 2022; Metzger 2022; Tuna and Metzger 2023; Wei et al. 2024). Furthermore, dust formation by the jets launched from the CBD could result in detectable IR radiation (see Grichener 2025 for a discussion on potential observational signatures of such systems). In the case of much wider post-asymptotic giant branch binaries, having orbital periods of $\simeq 100-1000$ days and masses of $M_{1}\simeq 1M_{\odot}$, mass accretion and jets launching from a CBD of several AU could shape a bipolar nebula (e.g., Bollen et al. 2022; Verhamme et al. 2024). These CBDs, however, are expected to be post-grazing envelope evolution disks rather than CEE (e.g., Soker 2020).
@@
One approach to allow a wide exploration of CEE and post-CEE outcomes is parameterizing processes. This is particularly critical in performing population synthesis studies. We aim to incorporate the possibility of forming a post-CEE CBD into population synthesis; here, we re-analyze the data from Grichener (2023) (publicly available at Zenodo: doi:10.5281) generated by the rapid population synthesis code compas (Riley et al., 2022), as described in Section II. Towards our goal, in Section III, we build a toy model to parameterize the conditions to form post-CEE CBDs. In Section IV we employ our toy model to the population synthesis data of massive binary systems that end with either NS - NS, BH - BH, or NS - BH binary systems. In Section V, we discuss the potential implications of our results for the formation and merger rates of double compact object (DCO) binaries. We summarize and discuss the effect of extending the parameter space on our results in Section VI.
II Methods
@@
We use rapid population synthesis to follow the evolution of massive binaries and find the properties of systems where a circumbinary disk would form in the post-CEE phase of a compact object (NS or BH) and a giant star. We analyze datasets from Grichener (2023) [^1] Publicly available at Zenodo: doi:10.5281. generated by version 02.31.06 of the population synthesis code compas (Compact Object Mergers: Population Astrophysics and Statistics; Stevenson et al. 2017; Vigna-Gómez et al. 2018; Riley et al. 2022).
@@
Given initial distributions and parametric assumptions, compas evolves isolated binary systems using the analytical fits from Hurley et al. (2000) and Hurley et al. (2002). The mass distribution of the primary (initially more massive) star is determined by the Kroupa initial mass function in the form $dN/dM_{\rm 1,ZAMS}\propto M_{\rm 1,ZAMS}^{-2.3}$ (Kroupa, 2001). The range of the masses $5\leq M_{\rm 1,ZAMS}/M_{\odot}\leq 100$ was chosen to account for primaries that could result in a supernova explosion and the formation of a compact object. [^2] The stars with the lowest masses in our distribution can gain mass through mass transfer and become massive enough to explode as supernovae. The mass of the secondary (initially less massive) star was determined by a flat distribution for a mass ratio between $0.1\leq q_{\text{ZAMS}}\equiv M_{2,\text{ZAMS}}/\penalty 50M_{1,\text{ZAMS}}\leq 1$ and the semi-major axis was determined by a flat-in-log distribution in the range $0.1\leq a_{\rm ZAMS}\rm/AU\leq 1000$ (Sana et al., 2012). For the NS natal-kicks induced by supernova explosions we take a bimodal velocity distribution with a higher mode of $\sigma_{\rm high}=265~\rm{km}~\rm{s}^{-1}$ for regular core collapse supernovae and a lower mode of $\sigma_{\rm high}=30~\rm{km}~\rm{s}^{-1}$ corresponding to electron capture supernovae and ultra stripped supernovae, while for the BHs in our sample we reduce this natal-kick distribution by a fallback factor. We focus on sub-populations of $10^{7}$ binaries with solar metallicity ($Z_{\rm\odot}=0.0142$; Asplund et al. 2009) and common envelope efficiency parameter $\alpha_{\rm CE}=1$ (equation 1) as in the ’fiducial’ model of Vigna-Gómez et al. (2018).
@@
In some of the systems sampled in Grichener (2023) there were episodes of dynamically unstable mass transfer that led to CEE. compas determines the stability of mass transfer by comparing the donor’s response to mass loss with that of the Roche-lobe radius: if the donor’s radius response, $d\ln R_{d,MT}/\penalty 50d\ln M_{d,MT}$, where $R_{d,MT}$ and $M_{d,MT}$ are the radius and mass of the donor at the onset of mass transfer, is greater than or equal to that of the Roche lobe, $d\ln R_{\mathrm{RL}}/\penalty 50d\ln M_{d,MT}$, where $R_{\mathrm{RL}}$ is the Roche radius, the mass transfer is stable. Otherwise, it results in a CEE event (e.g., Paczyński and Sienkiewicz 1972). To find the value of the semi-major axis immediately after the CEE phase, compas uses the energy formalism  [^3] Other methods use angular momentum parametrization to estimate the final orbital separation (e.g., Nelemans and Tout 2005; Toonen et al. 2012; Di Stefano et al. 2023). However, the angular momentum prescriptions could be problematic (e.g., Webbink 2008; Cohen and Soker 2023).. In this formalism, we compare the binding energy of the envelope $E_{\rm bind}$ to the orbital energy before and after the CEE phase (e.g., van den Heuvel 1976; Iben and Tutukov 1984; Livio and Soker 1988; Ivanova et al. 2013; Iaconi and De Marco 2019)
@@
where $\alpha_{\rm CE}$ is the common envelope efficiency parameter that determines which fraction of the orbital energy will go to unbinding the envelope, $M_{\rm 1,i}$ and $M_{\rm 2,i}$ are the masses of the primary and secondary stars at the onset of CEE, respectively, $M_{\rm 1,f}$ and $M_{\rm 2,f}$ are the masses of the primary and secondary stars after CEE, respectively, $a_{\rm i}$ is the semi-major axis of the binary systems before the CEE phase and $a_{\rm f}$ is the semi-major axis of the binary systems at the end of CEE. The binding energy is computed following de Kool (1990)
$E_{\rm bind}=-\frac{GM_{\rm i}M_{\rm env,i}}{\lambda R_{\rm i}},$ (2)
where $M_{\rm i}$, $M_{\rm env,i}$ and $R_{\rm i}$ are the mass, envelope mass and radius of the donor, respectively, and $\lambda$ is a structure parameter calculated by the prescription in Xu and Li (2010b) and Xu and Li (2010a). The structure parameter includes contributions of the full internal energy, which consists of the thermal energy, radiation energy, ionization energy and dissociation energy of molecular hydrogen. For more information about the population synthesis model assumptions and initial distributions see Section 2 in Grichener (2023).
@@
We focus on sub-populations in which a compact object enters a CEE phase with a giant star, where the formation of a CBD after CEE could potentially lead to shrinkage in the orbital separation and a post-CEE merger (e.g., Tuna and Metzger 2023; Wei et al. 2024), accounting for various luminous transients (e.g., Grichener 2025). Without taking the formation and interaction of the binary with the CBD into consideration, most of these systems end as DCO binaries, implying a potential reduction in their formation and merger rates (see Section V). compas allows us to examine which parameters and stellar quantities can affect the process of creating a CBD. To find the properties of the systems that will form a post-common envelope CBD, we compare the specific angular momentum of the envelope at the end of the Roche Lobe Overflow (RLOF) phase to the Keplerian specific angular momentum around the binary system, as we further elaborate in Section III.
III Parametrization of post-common envelope circumbinary disk formation
III.1 The toy model
@@
We build a toy model to parameterize the conditions for forming a post-CEE CBD and incorporate it in population synthesis studies. We estimate the fraction of the angular momentum at the onset of RLOF that the leftover envelope retains and equate it to the minimum angular momentum that the leftover requires to form a CBD. We then implement this parametrization in the data of Grichener (2023) and present our results in Sections IV and V.
@@
We assume that at the entrance to RLOF the secondary star, an NS or a BH in our study, manages to spin up the giant’s envelope to synchronization (e.g., Vigna-Gómez et al. 2020). We take the envelope to have a solid-body rotation with the synchronization angular velocity $\omega=\sqrt{G(M_{\rm*,RL}+M_{\rm CO,RL})/a_{\rm RL}^{3}}$, which corresponds to the Keplerian velocity of the compact object at the Roche-radius, where $M_{\rm*,RL}$ and $M_{\rm CO,RL}$ are the mass of the giant and the mass of the compact object at this stage, respectively, and $a_{\rm RL}$ is the orbital separation. For the moment of inertia of the envelope at this stage, we take $I_{\rm env,RL}=k_{\rm e}M_{\rm env,RL}R^{2}_{\rm\ast,RL}$, where $M_{\rm env,RL}$ and $R_{*,RL}$ are the envelope mass and radius of the giant at the onset of RLOF, and $k_{\rm e}\simeq 0.2$ for extended envelopes of giants (e.g., Soker 2004). The total angular momentum of the binary at the onset of RLOF is
$\begin{split}J_{\rm total,RL}=J_{\rm orb,RL}+J_{\rm rot,RL}=\\ \mu_{\rm RL}\sqrt{G(M_{\rm*,RL}+M_{\rm CO,RL})a_{\rm RL}(1-e^{2})}+\omega I_{\rm env,RL},\end{split}$ (3)
where $j_{\rm orb,RL}$ is the orbital angular momentum, $j_{\rm rot,RL}$ is the rotational angular momentum of the envelope and $\mu\equiv M_{\rm*,RL}M_{\rm CO,RL}/(M_{\rm*,RL}+M_{\rm CO,RL})$ at this stage, and $e$ is the eccentricity of the orbit [^4] In these systems we take $e$ prior to circularization at the onset of RLOF assumed in compas, which is often considerable due to the first SN’s natal kick. . We exclude the rotational angular momentum of the core and of the compact object, which are negligible compared to the rotational angular momentum of the envelope due to their extremely small radii.
Let us denote by $f_{\rm J}$ the fraction of angular momentum retained by the binary system after the CEE phase, and by $f_{\rm M}$ the fraction of mass that remains bound. We take the post-CEE specific angular momentum available to form a CBD to be a fraction $\beta\equiv f_{\rm J}/f_{\rm M}$ of the total specific angular momentum of the binary at the onset of RLOF (similarly to Izzard and Jermyn 2023)
$\begin{split}j_{\rm CBD}=\frac{f_{\rm J}J_{\rm total,RL}}{f_{\rm M}M_{\rm env,RL}}=\beta\frac{J_{\rm total,RL}}{M_{\rm env,RL}},\end{split}$ (4)
Post-CEE CBDs typically form with inner radii $R_{\rm CBD,in}$ approximately 2–3 times larger than the orbital separation of the binary at the end of the CEE phase (e.g., Kashi and Soker 2011; Tuna and Metzger 2023). The absolute minimal specific angular momentum required to form a thin ring that would viscously spread into a centrifugally-supported CBD is set by $R_{\rm CBD,in}$ . However, unlike the classic Lynden-Bell and Pringle (1974) disk description, which is less likely to form from CEE dynamics, we are interested in a disk which already extends from the inner stable orbit to a much larger radius, and thus has significantly larger angular momentum than a thin ring of the same mass at the inner stable orbit. Therefore, we must allow for additional angular momentum to support the gas that is further out. This more restrictive approach is motivated by the structure found in the hydrodynamical simulation of Vetter et al. (2024), and is conceptually similar to assuming the circularization of the CBD gas at a larger effective radius. Our criterion for CBD formation is then
$j_{\rm CBD}\gtrsim\zeta\sqrt{G(M_{\rm core}+M_{\rm CO,f})R_{\rm CBD,in}},$ (5)
where $M_{\rm core}$ is the mass of the giant’s core that becomes stripped after CEE, $M_{\rm CO,f}$ is the mass of the compact object companion at this stage, $R_{\rm CBD,in}=2.5a_{\rm f}$ and $\zeta$ is a dimensionless parameter that accounts for the mass distribution of the disk. For a disk whose surface density is a simple power law $\Sigma(r)=\Sigma_{\rm 0}(r/r_{\rm 0})^{-m}$, where $\Sigma_{\rm 0}$ and $r_{\rm 0}$ are normalization constants, calculating the mean specific angular momentum gives $\zeta=\frac{m-2}{m-2.5}$ (see Appendix for the derivation). The hydrodynamical simulations performed by Vetter et al. (2024) yield a more complicated structure. In particular, their disk has a clear outer radius, beyond which the gas is flowing radially outward rather than around the binary. We take a fiducial value of $\zeta=4$, which corresponds to $m\simeq 2.75$ and is roughly consistent with the density profile obtained by Vetter et al. (2024).Setting $\zeta=4$ can equivalently be interpreted as assuming that the gas circularizes into a CBD with an inner radius $R_{\rm CBD,in}$ that is 32 to 48 times the post CEE orbital separation. Such values are reasonable in systems with a large leftover envelope. Smaller values of $\beta$ will result in more DCOs that form post CEE CBDs. In addition, we assume that the mass of the disk is much smaller than the total mass of the binary system, and hence we can neglect its self gravity.
III.2 The role of the parameters $\beta$ and $\zeta$